Consider a marshmallow box containing 13 white, 5 orange, and 10 blue marshmallows.

1. We pick randomly one piece. What is the probability that we get
1. A blue marshmallow
2. A white marshmallow 3. An orang marshmallow


2. We pick 2 marshmallows. What is the probability that we get

1.2 marshmallows of the same color
2. at least one red marshmallow
1
Expert's answer
2020-12-25T15:24:17-0500

"\text{the probability that we get} \\
 \text{1.1. A blue marshmallow} =\frac{10}{28}\\
1.2.\text{ A white marshmallow}=\frac{13}{28}\\
1.3. \text{An orang marshmallow} =\frac{5}{28}\\
2.1. \text{marshmallows of the same color} \\=\frac{10}{28}\times\frac{9}{27}+\frac{13}{28}\times\frac{12}{27}+\frac{5}{28}\times\frac{4}{27}\\
=\frac{266}{756}\\
2.2. \text{ at least one red marshmallow}=0"